OpenStudy (anonymous):

slope:0.8, y-intercept:10

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

that is the question... here are the options..

OpenStudy (anonymous):

y=-0.8x+10 y=0.8x-10 y=0.8+10 y=5/7+10

OpenStudy (anonymous):

ok for this problem use the slope intercept equation which states: y-y1=m(x+x1), where m is the slope and (x1,y1) correspond to a coordinate. . Our coordinate for this problem is (0,10). Our slope is 0.8.(m). For the preliminary equation look like this: y-10=0.8(x+0). Then we simplify: y-10=0.8x+0, now add 10 to both sides of the equation. We get: y=0.8x+10

OpenStudy (anonymous):

My bad, the equations is y-y1=m(x-x1), but we still get y=0.8x+10

OpenStudy (anonymous):

hey!! thanks!!